# AoI-based Scheduling of Correlated Sources for Timely Inference

### v1

@@ V-C Performance Analysis
In this section, we analyze the performance of Algorithm 1 with an optimal policy in an asymptotic region. Let $\pi_{\text{gain}}$ denote the policy presented in Algorithm 1. Following the standard practice, we consider a set of bandits to be in the same class if they share identical penalty functions and transition probabilities.
Consider a base system with $M$ bandits, $N$ channels. Let $\mathrm{V}^{r}_{\mathrm{gain}}(\lambda)$ represent the objective value under policy $\pi_{\text{gain}}$ associated with Langrange multiplier $\lambda$ for a system that includes $rM$ bandits and $rN$ channels. We denote the fraction of class $m$ bandits in state $\delta$ at time $t$ by $\Phi^{m}_{\delta}(t)$. Then, we define
$\phi^{m}_{\delta}=\limsup_{T\rightarrow\infty}\sum_{t=0}^{T-1}\frac{1}{T}\mathbb{E}[\Phi^{m}_{\delta}(t)].$ (32)
We further use the vectors $\mathbf{\Phi}^{m}(t)$ and $\mathbf{\phi}^{m}$ to contain $\Phi^{m}_{\delta}(t)$ and $\phi^{m}_{\delta}$, respectively for all $\delta=1,2,\ldots,\delta_{\mathrm{bound}}$. Truncated AoI space will have little to no impact if $\delta_{\mathrm{bound}}$ is very large. This is because inference error achieves a stationary point as the AoI $\delta$ increases to a large value.
For a policy $\pi$, we can have the following mapping
$\Psi_{\pi}((\mathbf{\psi}^{m})_{m=1}^{M})=$
$\mathbb{E}[(\mathbf{\Phi}^{n}(t+1))_{n=1}^{N}|(\mathbf{\Phi}^{n}(t))_{n=1}^{N}=(\mathbf{\phi}^{n})_{n=1}^{N}].$ (33)
We define the $t$-th iteration of the maps $\Psi_{\pi,t\geq 0}(\cdot)$ as follows
$\Psi_{\pi,0}((\mathbf{\phi}^{m})_{m=1}^{M})=(\mathbf{\phi}^{m})_{m=1}^{M},$ (34)
$\Psi_{\pi,t+1}((\mathbf{\phi}^{m})_{m=1}^{M})=\Psi_{\pi}(\Psi_{\pi,t}((\mathbf{\phi}^{m})_{m=1}^{M})).$ (35)
Definition 3.
Uniform Global attractor. An equilibrium point $({\mathbf{\phi}^{m}_{\text{opt}}})_{m=1}^{M}$ given by the optimal solution of (V-B1) is a uniform global attractor of $\Psi_{\pi,t\geq 0}(\cdot)$, i.e., for all $\epsilon>0$, there exists $T(\epsilon)$ such that for all $t\geq T(\epsilon)$ and for all $({\mathbf{\phi}^{m}_{\text{opt}}})_{m=1}^{M})$, one has $||\Psi_{\pi,t}(({\mathbf{\phi}^{m}})_{m=1}^{M})-({\mathbf{\phi}^{m}_{\text{opt}}})_{m=1}^{M}))||_{1}\leq\epsilon$.
@@
If the Uniform Global attractor condition (Definition 3) holds and the number of sources $M$ and channels NN increase to $\infty$ while keeping the ratio $M/N$ constant, then we have
$\lim_{r\to\infty}|\mathrm{V}^{r}_{\mathrm{gain}}(\lambda^{*})-\mathrm{V}^{r}_{opt}|\leq\lim_{r\to\infty}\frac{2}{1-\gamma}\sum_{m=1}^{rM}O\bigg{(}\max_{n\neq m}\epsilon_{m,n}^{2}\bigg{)},$ (36)
where $\mathrm{V}^{r}_{opt}$ is the optimal objective value to (13)-(14) with rMrM bandits and $rN$ channels.
@@
See Appendix E. ∎
Theorem 3 tells us that if the correlation related term $\sum_{m=1}^{rM}\max_{m\neq n}\epsilon_{m,n}^{2}$ is close to zero, then our policy is optimal in an asymptotic region.
1 Input $\lambda_{1}=0$
@@
3  Evaluate $\alpha_{m,\lambda_{k}}$ for all $m$ and $\delta$
4   Update $\mathrm{Subgrad}\leftarrow 0$
@@ 7    Initialize $\pi_{m}(t)\leftarrow 0$ for all $m$
8    $\alpha_{m}\leftarrow\alpha_{m,\lambda_{k}}(\Delta_{m}(t))$
@@ 9      Schedule NN sources with highest $\alpha_{m}$
11 Update $\lambda_{k+1}\leftarrow\lambda_{k}+(\theta/k)(\mathrm{Subgrad}-\frac{N}{1-\gamma})$
@@ VI Design of Online Scheduling Policy with Unknown Penalty Function
  • We consider a bandit feedback structure: Receiver can compute the loss function $L(Y_{m,t},a_{m,t})$ associated with source $m$ only when it sends an update. Specifically, if at time slot $j$, source $m$ is scheduled, then $L(Y_{m,j},a_{m,j})$ can be computed, in which $Y_{m,j}$ can be inferred directly from the observation $X_{m,j}$. Receiver stores the loss value $L(Y_{m,j},a_{m,j})$ and the AoI $\Delta_{m}(j)$. At the end of every episode $k$, the receiver sends the stored loss values, AoI values, and the associated time-instants to the scheduler.
@@ VI-A Online Scheduling for Lagrangian Relaxed Problem
However, the estimate $\hat{f}_{m,k-1}(\delta)$ will naturally deviate from the true value. To account for this uncertainty, we define a confidence radius, $d_{m,k}(\delta)$ is the confidence radius with $0<\eta\leq 1$:
@@
which shrinks as we gather more samples for that state. This radius creates a ball of plausible values, $B_{m,k}(\delta)$:
@@
This optimistic value is then used in a value iteration algorithm to update our policy, ensuring we explore efficiently while accounting for statistical uncertainty:
@@
where $J^{(m)}(\delta)$ can obtained by using value iteration algorithm [46]. Then, we use the following estimated value function for all $\delta$:
@@ 9          $\pi_{m}(t)\leftarrow 1$; Schedule Source $m$
1 Input: Lagrange Multiplier $\lambda_{1}=0$
@@ 8      Schedule NN sources with highest $\alpha_{m}$
10 Update $\lambda_{k+1}\leftarrow\lambda_{k}+(\theta/k)(\mathrm{Subgrad}-\frac{N}{1-\gamma})$
@@
(b) For any episode $k$ with probability $1-O(\eta)$, we have
@@ VII Simulation Results
In this section, we will illustrate the performance of our Algorithm 1 and Algorithm 3.
@@ VII-A Simulation Model
We use a model from [9] for designing the correlation structure between sources. At the beginning of every time-slot, each source $m$ collects information about its own state. In addition, with probability $p_{m,n}$, the update collected by source $m$ also contains information about the current state of source $n$. Motivating examples of such correlated sources are cameras with overlapping fields of view and sensors with spatial correlation between the processes being monitored.
@@ VIII Conclusion and Future Work
Appendix A Proof of Theorem 1

### v2

@@ V-C Performance Analysis
In this section, we analyze the performance of Algorithm 1 with an optimal policy.
Let $\pi_{\text{gain}}$ denote the policy presented in Algorithm 1 and $\mathrm{V}_{\mathrm{gain}}(\boldsymbol{\lambda},T)$ represent the objective value under policy $\pi_{\text{gain}}$ associated with Lagrange multiplier $\boldsymbol{\lambda}$.
@@
Given a discount factor $\gamma\in(0.5,1)$, an optimal tie-breaker function $\psi$, and the fixed ratio $r=N/M$, if $\boldsymbol{\lambda_{k}}=\boldsymbol{\lambda^{*}}$, then in episode $k$, we have
$\frac{|\mathrm{V}_{\mathrm{gain}}(\boldsymbol{\lambda^{*},T})-\mathrm{V}_{opt}(T)|}{M}$
$\leq\frac{2(1-\gamma^{T})}{(1-\gamma)}\frac{\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg)}{M}+O(\beta\sqrt{\frac{(1-r)r}{M}}),$ (30)
where
$\beta=\frac{2(2\gamma)^{T}}{2\gamma-1}.$
@@
See Appendix D. ∎
Theorem 2 shows the normalized optimality gap of our policy depends on the correlation parameter $\epsilon_{m,n}$ and the system size $M$. In general, for large system size i.e., $M\to\infty$, the optimality gap can be finite, specifically we have
$\lim_{M\to\infty}\frac{|\mathrm{V}_{\mathrm{gain}}(\boldsymbol{\lambda^{*},T})-\mathrm{V}_{opt}(T)|}{M}$
$\leq\lim_{M\to\infty}\frac{2(1-\gamma^{T})}{(1-\gamma)}\frac{\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg)}{M}.$ (31)
However, for a special case, if the sum of the correlation parameters satisfies
$\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg)=o(M),$ (32)
we have
$\lim_{M\to\infty}\frac{2(1-\gamma^{T})}{(1-\gamma)}\frac{\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg)}{M}=0.$ (33)
For distributed sensor networks with a large number of sensors, where only a few sensors exhibit mutual correlation, the special case (32) can hold.
V-D Algorithm Simplification
In Algorithm 1, we consider a large sequence of Lagrange multiplier $\boldsymbol{\lambda}=(\lambda_{0},\lambda_{2},\ldots,\lambda_{T-1})$. This leads to a Lagrangian dual problem (22) that is practically difficult to solve in optimality. To address this challenge, we can simplify Algorithm 1 by considering $\lambda_{t}=\lambda$ for all $t$. For this modification, our Lagrange update rule becomes
$\lambda_{k+1}=$ $\lambda_{k}+\frac{\theta}{k}\bigg(\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}\mathbf{1}\bigg(\alpha_{m,\lambda_{k}}(\Delta_{m}(t))>\frac{\lambda_{k}}{\gamma}\bigg)$
$-\frac{(1-\gamma^{T})N}{1-\gamma}\bigg).$ (34)
1Input $\boldsymbol{\lambda_{k}}=0$
@@ V-D Algorithm Simplification
3  Evaluate $\alpha_{m,\boldsymbol{\lambda_{t,k},t}}(\delta)$ for all $m$ and $\delta$
4 Update $\mathrm{Subgrad}(t)\leftarrow 0$
@@ 7    Initialize $\pi_{m}(t)\leftarrow 0$ for all $m$
8    $\alpha_{m}\leftarrow\alpha_{m,\boldsymbol{\lambda_{t,k},t}}(\Delta_{m}(t))$
@@ 9      Schedule NN sources with highest $\alpha_{m}$
10     Sources with same gain index are ranked
11     with a tie-breaker index function $\psi_{t,m}$
12    $\mathrm{Subgrad}(t)\leftarrow\sum_{m=1}^{M}\gamma^{t}\mathbf{1}(\alpha_{m}>\frac{\lambda_{t,k}}{\gamma})-N$
13 Update $\boldsymbol{\lambda_{k+1}}$ using (V-B2)
@@ VI Design of Online Scheduling Policy with Unknown Penalty Function
  • Computing the loss function for all time $t$ and in real-time is computationally infeasible because ground truth target is not immediately available. To address this challenge, we adopt a bandit feedback structure. This structure allows the receiver to compute the loss function only for the scheduled time slots. The receiver stores the following quantities for every source $m$: the predicted output $a_{m,t}$, the most recent observation $X_{m,t-\Delta_{m}(t)}$, and $\pi_{m}(t)$ or equivalently $\Delta_{m}(t)$. Let’s consider $\pi_{m}(j)=1$, i.e., source $m$ is scheduled at time $t=j$. In this case, the receiver obtains $X_{m,j}$ at time $t=j+1$. Since we have assumed the true target $Y_{m,t}$ can be inferred directly from $X_{m,t}$ (e.g., $Y_{m,t}$ can be a function of, or equivalent to, $X_{m,t}$), the receiver can use the stored $X_{m,j}$ to compute $Y_{m,j}$. With both the prediction $a_{m,j}$ and the computed true label $Y_{m,j}$, the receiver can exactly compute the loss $L(Y_{m,j},a_{m,j})$. Crucially, this loss is only available at the scheduled time slots $j$ and not for all $t$, which defines the partial feedback characteristic of the bandit setting.
@@ VI-A Online Scheduling for Lagrangian Relaxed Problem
However, the estimate $\hat{f}_{m,k-1}(\delta)$ will naturally deviate from the true value. Given a chosen exploration parameter $\eta\in(0,1)$, we account for the uncertainty by defining the following confidence radius
@@
which shrinks as we gather more samples for that state. Moreover, the parameter $\eta$ also controls the confidence radius: The confidence radius increases as the parameter $\eta$ decreases.
The confidence radius creates a ball of plausible values, $B_{m,k}(\delta)$:
@@
Lemma 4 shows that the actual value $f_{m,k}(\delta)$ is in the ball of plausible values, $B_{m,k}(\delta)$ with probability $1-\eta$. As the parameter $\eta$ decreases, the ball $B_{m,k}(\delta)$ expands and the uncertainty reduces. However, as we will see next that the optimistic estimate of the actual value deviates away from the empirical estimate as $\eta$ decreases.
@@
As the parameter $\eta$ decreases, the confidence radius $d_{m,k}(\delta)$ increases and the estimate $\tilde{f}_{m,{k-1}}(\delta)$ deviates away from empirical value.
This optimistic value $\tilde{f}_{m,{k-1}}(\delta)$ is used in a value iteration algorithm to update our policy, ensuring we explore efficiently while accounting for statistical uncertainty:
@@
where $J^{(m)}(\delta)$ can obtained by using value iteration algorithm [45].
To ensure convergence of the estimated value function, we introduce a weighting parameter $\zeta\in(0,1)$. This parameter controls the trade-off between relying on the optimistic value function and retaining information from the history. Using $\zeta$, the estimated value function is defined as follows for all $\delta$:
@@ 9          $\pi_{m}(t)\leftarrow 1$; Schedule Source $m$
1Input: Lagrange Multiplier $\lambda_{1}=0$
@@ 8      Schedule NN sources with highest $\alpha_{m}$
Update $\lambda_{k+1}\leftarrow\lambda_{k}+(\theta/k)(\mathrm{Subgrad}-\frac{N}{1-\gamma})$
@@
(b) Given an exploration parameter $\eta\in(0,1)$, for any episode $k$ with probability $1-\eta$, we have
@@ VII Simulation Results
Though our theoretical analysis consider finite space $\mathcal{X}_{m}$ and $\mathcal{Y}_{m}$, our algorithms are general and can be applied to continuous space. Our algorithms only require countable AoI values. In this section, we will illustrate the performance of our Algorithm 1 and Algorithm 3 on a continuous space $\mathcal{X}_{m}$ and $\mathcal{Y}_{m}$ setting.
@@ VII-A Simulation Model
We use a model from [9] for designing the correlation structure between sources. At the beginning of every time-slot, each source $m$ collects information about its own state. In addition, with probability $p_{m,n}$, the update collected by source $m$ also contains information about the current state of source $n$. Motivating examples of such correlated sources are cameras with overlapping fields of view and sensors with spatial correlation between the processes being monitored. Here are two examples:
Cameras with Overlapping Fields of View: If camera $m$ successfully updates, there is a probability ($p_{m,n}$) that it also provides an updated view of targets associated with an adjacent camera $n$.
Spatially Correlated Environmental Sensors: Sensors monitoring processes (e.g., temperature, air quality, fire, rain) often exhibit spatial correlation. The probability ($p_{m,n}$) that one sensor’s update contains useful, non-redundant information about an adjacent sensor’s state.
@@ VIII Conclusion and Future Work
Appendix A How to Compute Penalty Function $f_{m}(\delta)$
Example 1.
$X_{m,t}$ $=a_{m}X_{m,t-1}+W_{m,t},$ (51)
[…]
where
$\bar{q}_{m,m}=q_{m,m}-\mathbf{q}_{m}Q^{-1}_{-1}\mathbf{q}_{m}^{T}$
$\mathbf{q}_{m}=[q_{m,1},\ldots,q_{m,m-1},q_{m,m+1},\ldots,q_{m,M}]$
Example 2.
$f_{m}(\delta)=\frac{\sum_{t=\delta+1}^{n}L(Y_{m,t},\phi_{m}(X_{m,t-\delta},Z_{-m,t-1}))}{n-\delta}.$ (53)
Appendix B Proof of Lemma 2
For the simplicity of presentation, we prove Lemma 2 for $M=3$ and $m=1$.
Part (a): we have
$g_{1}(\delta_{1},\delta_{2},\delta_{3})$ $=H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$\geq H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-\delta_{3}})$
$\geq H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-1})$
$=f_{m}(\delta_{1}),$ (54)
where the inequalities hold due to [43, Lemma 12].
Part (b): We have
$g_{1}(\delta_{1},\delta_{2},\delta_{3})$ $=H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$=H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$\quad-H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}},X_{2,t-1})$
$+H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$\overset{(c)}{=}H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-1})$
$+I_{L}(Y_{1,t};X_{3,t-1}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-\delta_{3}})$
$\quad+I_{L}(Y_{1,t};X_{2,t-1}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}}),$
$\overset{(d)}{\leq}f_{m}(\delta_{1})+\epsilon_{1,3}^{2}+\epsilon_{1,2}^{2}\leq f_{m}(\delta_{1})+2\max_{n\neq 1}\epsilon_{1,n}^{2}$
$\leq f_{m}(\delta_{1})+O(\max_{n\neq 1}\epsilon_{1,n}^{2}),$ (55)
where (a) holds due to Markov chain $Y_{1,t}\leftrightarrow X_{2,t-1}\leftrightarrow X_{2,t-\delta_{2}}$ [43, Lemma 12], (b) is obtained by definition (8), (c) can be obtained by using similar steps of (a)&(b), and (d) holds because of (2). This completes the proof for $M=3$ and $m=1$.
Following the same steps, we can prove it for any values of $M$ and $m$.
Appendix C Proof of Theorem 1
We define
Then, from (12)-(13) and (18)-(19), we have
$\mathrm{V}_{opt}(T)$
$\inf_{\pi\in\Pi_{c}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right]$
$\overset{a}{=}\inf_{\pi\in\Pi_{c}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right]$
$+\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}O\bigg(\max_{n=\neq m}\epsilon_{m,n}^{2}\bigg)$
$=\mathrm{V}_{f,opt}(T)+\frac{1-\gamma^{T}}{1-\gamma}\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg),$ (57)
where (a) is due to (16).
Moreover, we can write
$\mathrm{V}_{\pi_{f,opt}}(T)$
$=\mathbb{E}_{\pi_{f,opt}}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right]$
$\overset{b}{=}\mathbb{E}_{\pi_{f,opt}}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right]$
$+\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}O\bigg(\max_{n=\neq m}\epsilon_{m,n}^{2}\bigg)$
$=\mathrm{V}_{f,opt}(T)+\frac{1-\gamma^{T}}{1-\gamma}\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg),$ (58)
where (b) is due to (16).
By combining (C) and (C), we obtain (20).
Appendix D Proof of Theorem 2
By using triangle inequality, We can write
$|\mathrm{V}_{\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)-\mathrm{V}_{opt}(T)|$ $\leq|\mathrm{V}_{\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)-\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)|$
$+|\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)-V_{f,opt}(T)|$
$+|V_{f,opt}(T)-V_{opt}(T)|,$ (59)
where we define
$\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)=\mathbb{E}_{\pi_{gain}}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right].$ (60)
From (C), we obtain
$|V_{f,opt}(T)-V_{opt}(T)|=\frac{1-\gamma^{T}}{1-\gamma}\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg)$ (61)
Similar to (C), we can also get
$\mathrm{V}_{\mathrm{gain}}(\boldsymbol{\lambda^{*},T})$
$=\mathbb{E}_{\pi_{gain}}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right]$
${=}\mathbb{E}_{\pi_{gain}}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right]$
$+\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}O\bigg(\max_{n=\neq m}\epsilon_{m,n}^{2}\bigg)$
$=\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*},T})+\frac{1-\gamma^{T}}{1-\gamma}\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg).$ (62)
For an optimal tie-breaker function $\psi$ the gain index-based policy satisfies [47, Proposition 5]:
$\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)-V_{f,opt}(T)$ $\leq\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)-\mathrm{V}_{f,opt}(\boldsymbol{\lambda^{*}},T)$
$\leq\sum_{t=0}^{T-1}\beta_{t}\sqrt{N(1-r)},$ (63)
where $\beta_{t}=\frac{\gamma^{t}}{2\gamma-1}((2\gamma)^{T-t}-1)$ as provided in [47, equation (24)]. Next, we have [47, equation (59)]
$\sum_{t=0}^{T-1}\beta_{t}=\frac{1}{2\gamma-1}[2\gamma^{T}(2^{T}-1)-\frac{1-\gamma^{T}}{1-\gamma}].$ (64)
For $\delta>0.5$, we have ${2\gamma-1}>0$. Then, we can drop the negative terms
$\frac{2\gamma^{T}}{2\gamma-1}-\frac{1-\gamma^{T}}{1-\gamma}$
and get
$\mathrm{V}_{f,\mathrm{gain}}(\boldsymbol{\lambda^{*}},T)-V_{f,opt}(T)$
$\leq O(\beta\sqrt{N(1-r)})=O(\beta\sqrt{Mr(1-r)}),$ (65)
where $\beta=\frac{2(2\gamma)^{T}}{2\gamma-1}$.
By substituting (61)-(D) into (D), we obtain Theorem 2.
Appendix E Proof of Theorem 3
This proves Theorem 3(a).
Part (b) To prove Theorem 3(b), the following Lemma is important.
Lemma 4.
For the penalty function $f_{m,k}$, we have
$\mathrm{Pr}\bigg(f_{m,k}(\delta)\in B_{m,k}(\delta)\bigg)=1-\eta.$ (66)
Lemma 5 (Hoeffding’s inequality).
Let $Z_{1},Z_{2},\ldots,Z_{n}$ are i.i.d. samples of a random variable $Z\in[0,1].$ For any $\eta>0$, we must have
$\mathrm{Pr}\bigg\{\bigg|\mathbb{E}[Z]-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\bigg|\leq\sqrt{\frac{\mathrm{ln}(2/\bar{\eta})}{2n}}\bigg\}\geq 1-\eta.$ (67)
Now, using Lemma 4, (VI-A)-(VI-A), we prove Theorem 3(b). By using triangle inequality, we can have
$\bigg|J_{m,\lambda,{k}}(\delta)-J_{m,\lambda,{k-1}}(\delta)\bigg|$
$\leq|J_{m,\lambda,{k}}(\delta)-J_{m,\mathrm{optimistic},k}(\delta)|$
$+|J_{m,\mathrm{optimistic},k}(\delta)-J_{m,\lambda,{k-1}}(\delta)|.$ (68)
By using [52, Theorem 12], we can have with probability $1-\eta$:
$|\tilde{f}_{m,k-1}(\delta)-f_{m,k-1}(\delta)|\leq d_{m,k-1}(\delta).$ (69)
This yields:
$|J_{m,\mathrm{optimistic},k}(\delta)-J_{m,\lambda,{k-1}}(\delta)|$
${\leq}\frac{1}{1-\gamma}\max_{\delta^{\prime}}d_{m,k-1}(\delta^{\prime})$ (70)
and
$|J_{m,\lambda,{k}}(\delta)-J_{m,\mathrm{optimistic},k}(\delta)|$
$+\frac{1}{1-\gamma}\max_{\delta^{\prime}}d_{m,k-1}(\delta^{\prime}).$ (71)
By substituting (E), (E) into (E), we obtain Theorem 3(b).
Appendix F Minimization of Time-averaged Sum of Inference Errors
Our goal is to find a policy $\pi\in\Pi$ that minimizes the time-averaged sum of inference errors:
$\mathrm{L_{opt}}=$
$\inf_{\pi\in\Pi}\limsup_{T\to\inf}\mathbb{E}_{\pi}\frac{1}{T}\left[\sum_{t=0}^{T-1}\sum_{m=1}^{M}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right],$ (72)
$~\mathrm{s.t.}\sum_{m=1}^{M}\pi_{m}(t)=N,t=0,1,\ldots,$ (73)
where $g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))$ is the inference error for the $m$-th target at time $t$ and $\mathrm{L_{opt}}$ is the minimum average inference error.
For $M=2$ sources and $N=1$, we are able to find a low-complexity optimal policy to (F)-(73). According to Theorem 4, the optimal policy is a stationary cyclic policy.
Definition 2 (Stationary Cyclic Policy).
Theorem 4.
$+\sum_{j=0}^{\tau_{2}-1}\sum_{m=1}^{2}\big(g_{m}(2+j,1)\big)\bigg),$ (74)
where $\mathrm{L_{opt}}$ is the optimal objective value of (12)-(13).
Proof.